# Otter: Generating Tests from Issues to Validate SWE Patches

### v1

@@ Abstract
While there has been plenty of work on generating tests from existing code, there has been limited work on generating tests from issues. A correct test must validate the code patch that resolves the issue. In this work, we focus on the scenario where the code patch does not exist yet. This approach supports two major use-cases. First, it supports TDD (test-driven development), the discipline of “test first, write code later” that has well-documented benefits for human software engineers. Second, it also validates SWE (software engineering) agents, which generate code patches for resolving issues. This paper introduces Otter, an LLM-based solution for generating tests from issues. Otter augments LLMs with rule-based analysis to check and repair their outputs, and introduces a novel self-reflective action planning stage. Experiments show Otter outperforming state-of-the-art systems for generating tests from issues, in addition to enhancing systems that generate patches from issues. We hope that Otter helps make developers more productive at resolving issues and leads to more robust, well-tested code.
@@ 1 Introduction
A software engineering (SWE) issue is a bug report or feature request for improving the code in a repository. Before a software engineer attempts to write new code that resolves a bug, they typically write a reproduction test to confirm the presence of the bug in the old code. In fact, even when working on a feature request, before writing new code it is recommended to create an acceptance test first, to confirm the presence of the feature in the new code once written. This practice is known as test-driven development (TDD), and it improves the quality of both tests and the code itself (Beck, 2002). There has been recent work on resolving SWE issues automatically with LLMs, typically with agents (Yang et al., 2024; Ruan et al., 2024; Wang et al., 2024b). Since the release of SWE-Bench (Jimenez et al., 2024), such issue-resolution systems are often called SWE agents. We posit that, just like up-front tests help human SWE engineers, they also help automated SWE agents.
@@
This paper introduces Otter (acronym for “Otter: TDD-Test gEnerator for Reproducing issues”). It takes as input an issue description and the original code in the repository, and generates as output a set of tests . Otter contains a novel self-reflective action planner for deciding which code to read and which tests to write. It also uses rule-based code analyses and transformations throughout the workflow to curate LLM inputs, validate LLM outputs, and repair generated tests. Otter generates fail-to-pass tests in 31.4% of cases, and in an ensemble of size 5 dubbed Otter++, that increases to 37.0%. At the same time, the cost (even with ensembling) is less than $0.10 per issue with GPT-4o.
The most popular benchmark for automatically resolving SWE issues is SWE-bench Verified (Chowdhury et al., 2024). To evaluate solutions such as Otter that automatically generate tests from issues, we introduced a new benchmark, TDD-Bench-Verified (Ahmed et al., 2024) in our prior work. TDD-Bench-Verified 111https://github.com/IBM/TDD-Bench-Verified evaluates tests by checking whether the tests a) fail on the old code before issue resolution, b) pass on the new code, and c) cover the code changes well. The fact that TDD-Bench-Verified is derived from SWE-bench Verified enables us to empirically study the effects of generated tests on SWE agents. We observe that the tests from Otter++ can be used to trade precision for recall on the SWE-bench Verified leaderboard. E.g., for the system ranked 3rd on the leaderboard, filtering by the generated tests boosts precision from 60.8% to 91.9% while reducing recall to 33%. This paper makes the following contributions:
@@ 2 Problem Statement
In this work, we focus on the problem of generating tests from issues. Specifically, the input is a pair of an issue description alongside the old version of the code before the issue is resolved. The issue description is typically in natural language, but it may sometimes contain embedded code snippets or stack traces. The code is a snapshot of all the files and folders in a Python source code repository. The expected output is a set of tests that should go from failing on to passing on , which is the new version of the code after the issue is resolved. By failing on , the tests reproduce the issue, and by passing on , they validate its resolution. Besides going from failing to passing, the tests should also maximize coverage of the code change. A solution to this problem is thus a function that takes an input and returns tests . The new code is not available to , which must generate tests based on alone. This is representative of the real world where source code repositories may have regression tests for existing code but lack tests for open issues. Otter provides an implementation of the function and, thus, a solution to this problem.
@@ 3 Related Work
When resolving issues, some SWE agents also generate tests along the way. The original SWE-Agent (Yang et al., 2024), a single-agent system, attempts to reproduce the issue, as explicitly specified in its prompt. Some multi-agent systems—CodeR (Chen et al., 2024) and SpecRover (Ruan et al., 2024)—start with a Reproducer agent for generating tests. However, none of the three (SWE-Agent, CodeR, or SpecRover) are evaluated for the effectiveness of their generated tests. Agentless (Xia et al., 2024) relies on inference scaling, generating several candidate patches and several tests. It then uses the tests to help rank the patches, ultimately choosing a single patch to submit. The effectiveness of the tests is evaluated indirectly by their impact on issue resolution rate (from 27% to 32%), not directly for their own fail-to-pass rate or coverage like in our work.
Two very recent systems are dedicated to generating tests from Python issues. Aegis (Wang et al., 2024a) is a multi-agent system that uses inference scaling, but the exact dataset for their evaluation is unclear (the paper says SWE-bench Lite, but then compares against numbers from another system on SWT-bench Lite, which is different). Aegis is more costly than Otter++, and unlike our work, the Aegis paper (a) does not report coverage numbers and (b) does not evaluate how tests can help trade off precision vs. recall w.r.t. performance of SWE agents. SWE-Agent+ adapts a patch-generating agent to generate tests instead (Mündler et al., 2024) and achieves 19.2% fail-to-pass rate on the SWT-bench Lite dataset introduced by the same paper. SWT-bench applies less rigorous quality filters than TDD-Bench-Verified, and it measures coverage in a round-about way by first running additional tests than just the generated ones and then subtracting them back out. Using tests generated by SWE-Agent+ as a filter improves precision of SWE agents to 47.8% while reducing recall to 20%. Otter++ outperforms SWE-Agent+ on all of these metrics.
@@ 4.1 Otter: Test Generation Guided by Self-reflective Action Planner
Before developers tackle an issue, they usually start by gaining an understanding of the current state of the project. They do this by localizing relevant existing tests and focal functions (i.e., functions exercised by those tests and likely places for a fix). Depending on the developer’s familiarity with the project and the nature of the issue, the difficulty of localization may vary. Inspired by human developers’ actions, our approach also starts with test and focal function localization. The localizer phase collects all files from that contain at least one test function for test localization. It presents the list of files and the issue description to the LLM and asks it to generate 10 relevant test file names. Our initial findings on the dev set of SWE-bench indicated that the Top-10, Top-5, and Top-1 accuracy for test file localization with GPT-4o are 83.6%, 76.0%, and 59.1%, respectively. We restricted file retrieval to 10 files so as not to overwhelm the contexts for subsequent LLM calls. Next, the localizer validates the file names by comparing them with the previously collected file list and drops the ones that do not match. After localizing test files, it makes a second LLM call with file names and test function names from those files. The model chooses the test files and functions relevant to . It again validates the retrieved file names, but instead of dropping any hallucinated file names, replaces them with the file names with minimal edit distance. This ensures validity of identified files, which is essential for the subsequent test-generation phase. Note that even if the localizer chooses the wrong test file, if the file is at least valid, the test generator may still succeed. Focal file and function localization follows a similar two-LLM-call approach. Figures 13 and 13 in the Appendix illustrate the LLM prompts for localization.
@@
The second phase of Otter creates a plan, which is a list of actions for generating the fail-to-pass tests . There are three kinds of actions: read, write, or modify. A read action reads a function from to use as context in a prompt. A write or modify action declares the intent to write a new test function or modify an existing test function. Here, is a file and function name (in the planner, write and modify actions do not yet include the exact code for the test function, which is left to the test generator phase of Otter). The planner starts by executing read actions for the files and functions provided by the localizer. Next, it prompts an LLM with the function definitions, the issue description , and instructions to make an initial plan, comprising only read actions. The next step of the planner validates the planned actions: it checks whether the actions generated by the LLM refer to valid file and function names. The final step of the planner is “reflect and improve plan”, an LLM call with a prompt including feedback from validation. At this point, the plan is no longer restricted to only read actions, and can also contain a write or modify action. We provide statistics on read/write/modify actions generated by the Planner in Section B. The model is also instructed to self-reflect on the proposed plan with one of three possible outcomes: “Satisfied”, “Unsatisfied”, and “Unsure”. If the model chooses “Satisfied”, Otter moves forward to the test-generation phase. For other options, it returns to the validation step and then repeats the “reflect and improve plan” step. This process is repeated at most five times. In most cases, the model is satisfied with the plan in the first two turns. The two planner prompts are presented in Figures 17 and 17 (Appendix).
@@
The test-generation phase of Otter executes the actions computed by the planner, which can involve generating a new test (write action) or updating an existing test (modify action). To guide the LLM in this task, we extract the test structure and imports from the localized test file and make them available in the prompt. This reduces the burden on the LLM to generate imports. The file structure is relevant for new tests, to determine their insertion point in the test file. Otter uses a different prompt for a write vs. a modify action, illustrated in Figures 19 and 19 in the Appendix. For new tests, the model needs to generate the preceding function name in addition to the test. Unlike Mündler et al. (2024), we did not try to generate diffs; instead, we ask the model to generate complete test functions, even for the modification case. Since model training data tends to contain more complete functions than diffs, we expect the model to perform better at generating functions.
@@
4.2 Otter++: Ensemble Method using Multi-Sampling with Heterogeneous Prompting
Otter has several components (e.g., localizer, planner), but they are not perfect. The file localizer accuracy for focal and test localization is 82.4% and 70.6% with GPT-4o, respectively. As the output from localization serves as input to later LLM calls, those later calls may be affected by inaccuracies in localization. Conversely, LLMs can sometimes generate fail-to-pass tests in a zero-shot setup, even without any context. So, selectively including and excluding parts of localization may produce different fail-to-pass tests that are not generated by Otter. Otter++ uses the test generated by Otter (T1) and adds four new tests (T2–T5), obtained by skipping the planner stage, and including neither, one, or both of focal and test localization. In other words, Otter++ runs the test generator stage five times with different, heterogenous prompts. We favor heterogeneous prompts with greedy decoding over homogeneous prompts with higher model temperature because our initial experiments showed that the latter yields poorer tests and lower diversity.
To pick the best among the five candidate tests, we run the five tests on (recall that is not available yet) and analyze the execution logs. If a test passes on , we just discard that test because it violates the fail-to-pass criterion. We classify the remaining tests into three groups: assertion failure, other failure (the test runs but produces the wrong output), and error (the test does not run properly, e.g., because of wrong syntax or an exception in a fixture). We pick a test from the first non-empty group to maximize the chance that it failed for the right reason (i.e., it reproduces the bug described in the issue). If the selected group has multiple tests, we break the tie with a pre-determined ordering of the five prompts that favors tests from prompts with more or better information: Otter, followed by the prompts with both localizers, test localizer only, focal localizer only, and neither localizer, in that order.

### v2

@@ Abstract
While there has been plenty of work on generating tests from existing code, there has been limited work on generating tests from issues. A correct test must validate the code patch that resolves the issue. This paper focuses on the scenario where that code patch does not yet exist. Doing so supports two major use-cases. First, it supports TDD (test-driven development), the discipline of “test first, write code later” that has well-documented benefits for human software engineers. Second, it also validates SWE (software engineering) agents, which generate code patches for resolving issues. This paper introduces TDD-Bench-Verified, a benchmark for generating tests from issues, and Otter, an LLM-based solution for this task. Otter augments LLMs with rule-based analysis to check and repair their outputs, and introduces a novel self-reflective action planner. Experiments show Otter outperforming state-of-the-art systems for generating tests from issues, in addition to enhancing systems that generate patches from issues. We hope that Otter helps make developers more productive at resolving issues and leads to more robust, well-tested code.
@@ 1 Introduction
A software engineering (SWE) issue is a bug report or feature request for improving the code in a repository. Before a software engineer attempts to write new code that resolves a bug, they typically write a reproduction test to confirm the presence of the bug in the old code. In fact, even when working on a feature request, before writing new code it is recommended to create an acceptance test first, to confirm the presence of the feature in the new code once written. This practice is known as test-driven development (TDD), and it improves the quality of both tests and the code itself (Beck, 2002). The recent introduction of SWE-Bench (Jimenez et al., 2024) has spurred work on resolving SWE issues automatically with LLMs, typically with agents (Yang et al., 2024; Ruan et al., 2024; Wang et al., 2024b). Such issue-resolution systems are often called SWE agents. We posit that, just like up-front tests help human SWE engineers, they also help automated SWE agents.
@@
This paper introduces Otter (acronym for “Otter: TDD-Test gEnerator for Reproducing issues”). It takes as input an issue description and the original code in the repository, and generates as output a set of tests . Otter contains a novel self-reflective action planner for deciding which code to read and which tests to write. It also uses rule-based code analyses and transformations throughout the workflow to curate LLM inputs, validate LLM outputs, and repair generated tests. Otter generates fail-to-pass tests in 31.4% of cases, and in an ensemble of size 5 dubbed Otter++, that increases to 37.0%. At the same time, the cost (even with ensembling) is less than $0.10 per issue with GPT-4o. Tests generated by Otter adhere to the same testing framework as repository’s existing test suite, to which they can be added.
Today the most popular benchmark for automatically resolving SWE issues is SWE-bench Verified (Chowdhury et al., 2024). To evaluate solutions such as Otter that automatically generate tests from issues, we introduce a new benchmark, TDD-Bench-Verified. TDD-Bench-Verified evaluates tests by checking whether the tests a) fail on the old code before issue resolution, b) pass on the new code, and c) cover the code changes well. The fact that TDD-Bench-Verified is derived from SWE-bench Verified enables us to empirically study the effects of generated tests on SWE agents. We observe that the tests from Otter++ can be used to trade precision for recall on the SWE-bench Verified leaderboard. For instance, for the system ranked 3rd on the leaderboard, filtering by the generated tests boosts precision from 60.8% to 91.9% while reducing recall to 33%. This paper makes the following contributions:
@@
TDD-Bench-Verified 111https://github.com/IBM/TDD-Bench-Verified, a benchmark for evaluating tests generated from issues, including a high-quality dataset and a metric based on test results and coverage.
@@ 2 Problem Statement
This work focuses on the problem of generating tests from issues. Specifically, the input is a pair of an issue description alongside the old version of the code before the issue is resolved. The issue description is typically in natural language, but it may sometimes contain embedded code snippets or stack traces. The code is a snapshot of all the files and folders in a Python source code repository. The expected output is a set of tests that should go from failing on to passing on , which is the new version of the code after the issue is resolved. By failing on , the tests reproduce the issue, and by passing on , they validate its resolution. Besides going from failing to passing, the tests should also maximize coverage of the code change (formalized in Section 5). A solution to this problem is thus a function that takes an input and returns tests . The new code is not available to , which must generate tests based on alone. This is representative of the real world where source code repositories may have regression tests for existing code but lack tests for open issues. Otter provides an implementation of the function and, thus, a solution to this problem.
@@ 3 Related Work
When resolving issues, some SWE agents also generate tests along the way. The original SWE-Agent (Yang et al., 2024), a single-agent system, attempts to reproduce the issue, as explicitly instructed in its prompt. Some multi-agent systems—CodeR (Chen et al., 2024) and SpecRover (Ruan et al., 2024)—start with a Reproducer agent for generating tests. However, none of the three (SWE-Agent, CodeR, or SpecRover) are evaluated for the effectiveness of their generated tests. Agentless (Xia et al., 2024) relies on inference scaling, generating several candidate patches and several tests. It then uses the tests to help rank the patches, ultimately choosing a single patch to submit. The effectiveness of the tests is evaluated indirectly by their impact on issue resolution rate (from 27% to 32%), not directly for their own fail-to-pass rate or coverage like in our work.
Three very recent systems are dedicated to generating tests from Python issues. Aegis (Wang et al., 2024a) is a multi-agent system that uses inference scaling, but the exact dataset for their evaluation is unclear (the paper says SWE-bench Lite, but then compares against numbers from another system on SWT-bench Lite, which is different). Aegis is more costly than Otter++, and unlike our work, the Aegis paper (a) does not report coverage numbers and (b) does not evaluate how tests can help trade off precision vs. recall w.r.t. performance of SWE agents. EvoCoder (Lin et al., 2024) uses experiences from prior issues to help with the latest issue at hand, which is complementary to Otter. Unlike our work, the generated tests are not integrated with the existing CI pipeline. Furthermore, the experiments do not use execution-based metrics, making it hard to compare empirically. SWE-Agent+ adapts a patch-generating agent to generate tests instead (Mündler et al., 2024) and achieves 19.2% fail-to-pass rate on the SWT-bench Lite dataset introduced by the same paper. SWT-bench applies less rigorous quality filters than TDD-Bench-Verified, and measures coverage in a round-about way by first running additional tests than just generated ones and then subtracting them back out. Using tests from SWE-Agent+ as a filter improves precision of SWE agents to 47.8% while reducing recall to 20%. Otter++ outperforms SWE-Agent+ on all of these metrics.
CodeT (Chen et al., 2023) is one of the first approaches that leverage the same LLM to automatically generate both code samples and corresponding test cases. CodeT then executes generated code samples using the generated test cases and performs a dual execution agreement to choose the best solution and test. CodeT is not directly applicable to our current setup because we have multiple tests but no code patch. CodeMonkeys (Ehrlich et al., 2025) is another relevant work that iteratively and jointly improves a patch with a test, which is interesting and complementary to Otter. Unlike our work, their experiments do not evaluate the fail-to-pass rate nor the coverage of generated tests.
@@ 4.1 Otter: Test Generation Guided by Self-reflective Action Planner
Before developers tackle an issue, they usually start by gaining an understanding of the current state of the project. They do this by localizing relevant existing tests and focal functions (i.e., functions exercised by those tests and likely places for a fix). Depending on the developer’s familiarity with the project and the nature of the issue, the difficulty of localization may vary. Inspired by human developers’ actions, our approach also starts with test and focal function localization. The localizer phase collects all files from that contain at least one test function for test localization. It presents the list of files and the issue description to the LLM and asks it to generate 10 relevant test file names. Our initial findings on the dev set of SWE-bench indicated that the Top-10, Top-5, and Top-1 accuracy for test file localization with GPT-4o are 83.6%, 76.0%, and 59.1%, respectively. We restricted file retrieval to 10 files so as not to overwhelm the contexts for subsequent LLM calls. Next, the localizer validates the file names by comparing them with the previously collected file list and drops the ones that do not match. After localizing test files, it makes a second LLM call with file names and test function names from those files. The model chooses the test files and functions relevant to . The localizer again validates the retrieved file names, but instead of dropping any hallucinated file names, replaces them with the file names with minimal edit distance. This ensures validity of identified files, which is essential for the subsequent test-generation phase. Note that even if the localizer chooses the wrong test file, if the file is at least valid, the test generator may still succeed. Focal file and function localization follows a similar two-LLM-call approach. Figures 15 and 15 in the Appendix illustrate the LLM prompts for localization.
@@
The second phase of Otter creates a plan, which is a list of actions for generating the fail-to-pass tests . There are three kinds of actions: read, write, or modify. A read action reads a function from to use as context in a prompt. A write or modify action declares the intent to write a new test function or modify an existing test function. Here, is a file and function name (in the planner, write and modify actions do not yet include the exact code for the test function, which is left to the test generator phase of Otter). The planner starts by executing read actions for the files and functions provided by the localizer. Next, it prompts an LLM with the function definitions, the issue description , and instructions to make an initial plan, restricted to only read actions. The next step of the planner validates the planned actions: it checks whether the actions generated by the LLM refer to valid file and function names. The final step of the planner is “reflect and improve plan”, an LLM call with a prompt including feedback from validation. At this point, the plan is no longer restricted to only read actions, and can also contain a write or modify action. Section B in the appendix provides statistics on read/write/modify actions generated by the planner. The model is also instructed to self-reflect on the proposed plan with one of three possible outcomes: “Satisfied”, “Unsatisfied”, and “Unsure”. If the model chooses “Satisfied”, Otter moves forward to the test-generation phase. For other options, it returns to the validation step and then repeats the “reflect and improve plan” step. This process is repeated at most five times. In most cases, the model is satisfied with the plan in the first two turns. The two planner prompts are presented in Figures 19 and 19 (Appendix).
@@
The test-generation phase of Otter executes the actions computed by the planner, which can involve generating a new test (write action) or updating an existing test (modify action). To guide the LLM in this task, we extract the test structure and imports from the localized test file and make them available in the prompt. This reduces the burden on the LLM to generate imports. The file structure is relevant for new tests, to determine their insertion point in the test file. Otter uses a different prompt for a write vs. a modify action, illustrated in Figures 21 and 21 in the Appendix. For new tests, the model needs to generate the preceding function name in addition to the test. Unlike most SWE agents, Otter does not try to generate diffs; instead, it asks the model to generate complete test functions, even for the modification case. Since model training data tends to contain more complete functions than diffs, we expect the model to perform better at generating functions.
@@
4.2 Otter++: Ensemble using Multi-Sampling with Heterogeneous Prompting
Otter has several components (e.g., localizer, planner), but they are not perfect. The file localizer accuracy for focal and test localization is 82.4% and 70.6% with GPT-4o, respectively. As the output from localization serves as input to later LLM calls, those later calls may be affected by inaccuracies in localization. Conversely, LLMs can sometimes generate fail-to-pass tests in a zero-shot setup, even without any context. So, selectively including and excluding parts of localization may produce different fail-to-pass tests that are not generated by Otter. Otter++ uses the test generated by Otter (T1) and adds four new tests (T2–T5), obtained by skipping the planner stage, and including neither, one, or both of focal and test localization. In other words, Otter++ runs the test generator stage five times with different, heterogeneous prompts. We favor heterogeneous prompts with greedy decoding over homogeneous prompts with higher model temperature because our initial experiments showed that the latter yields poorer tests and lower diversity.
To pick a single test among the five candidate tests, we run the five tests on (recall that is not available yet) and analyze the execution logs. If a test passes on , we just discard that test because it violates the fail-to-pass criterion. We classify the remaining tests into three groups: assertion failure, other failure (the test runs but produces the wrong output), and error (the test does not run properly, e.g., because of wrong syntax or an exception in a fixture). We pick a test from the first non-empty group to maximize the chance that it failed for the right reason (i.e., it reproduces the bug described in the issue). If the selected group has multiple tests, we break the tie with a pre-determined ordering of the five prompts that favors tests from prompts with more or better information: Otter, followed by the prompts with both localizers, test localizer only, focal localizer only, and neither localizer, in that order.
@@ 4.3 Baseline: Zero-shot Test File Generation
5 TDD-Bench-Verified Benchmark
This section presents TDD-Bench-Verified, a new benchmark that supports evaluation of techniques for generating tests from an issue description and an old code version, without access to new code that resolves the issue (see problem statement in Section 2).
5.1 TDD-Bench-Verified Evaluation Harness
Figure 3 shows the harness for evaluating tests , which typically come from a solution, but the harness can also be applied on golden tests mined from a pull request (PR). The evaluation harness runs in a containerized environment. Starting at the top left, tests come in the form of a patch, which is applied (via git apply) on the old code . Next, the harness analyzes the resulting code to resolve the contributed test functions . The harness then executes while avoiding running other tests that occur in the same file but were not contributed in the patch. This yields test results, including coverage achieved on the old code. At least one of these tests should fail, indicating that the tests reproduce the issue at hand.
Moving to the bottom part of Figure 3, the code changes come from the golden code patch mined from the same PR, which is applied on to obtain the golden new code . The harness executes the tests again, this time on the new code, to obtain a second set of test results. This time, all tests should pass, to validate that the issue was indeed resolved. An example test patch is presented in Figure 9.
5.2 Dataset Filters
TDD-Bench-Verified builds upon prior work from SWE-bench (Jimenez et al., 2024) and SWE-Bench Verified (Chowdhury et al., 2024). SWE-bench uses filters to keep only those mined instances for which the set of golden tests contains at least some tests that fail on and pass on the golden new code from the same PR. SWE-Bench Verified is a subset of SWE-Bench, consisting of 500 instances further vetted by human annotators (Chowdhury et al., 2024). The annotators filtered out instances where the issue description was underspecified or the golden tests were overly specific, i.e., would reject some valid new code . They also removed some instances where tests failed due to environment problems.
In the same spirit, TDD-Bench-Verified applies more filters to obtain an even higher-quality dataset. In a nutshell, the filtering process applies the TDD-Bench-Verified evaluation harness (Figure 3) to the golden tests from the original PR. Specifically, substituting wherever occurs in Figure 3 checks whether the PR indeed contributed tests that went from failing to passing. We filter out any instance where the contributed tests do not satisfy that criterion. Although the human annotators of SWE-bench Verified were diligent, a few instances slipped past their filters, and we drop those for TDD-Bench-Verified.
Figure 2 illustrates the filtering process. Starting from the 500 instances of SWE-bench Verified, we first drop 7 instances whose environment we could not recreate. Next, we run the test harness on the golden tests . This filters out 44 additional instances because the tests do not have the expected fail-to-pass behavior (25 instances) or have zero line coverage on the golden code patch (19 instances). In the end, 449 high-quality instances remain across 12 repositories. Table 7 summarizes key statistics of TDD-Bench-Verified.
5.3 Evaluation Metric
Passing a test does not necessarily mean a patch is adequate to address the issue. Aleithan et al. (2024) reported that 31.1% of the passed code patches in SWE-Bench are suspicious due to weak test cases. To evaluate test adequacy, we compute the coverage of the submitted test patch. One key difference between SWE-Bench Verified and TDD-Bench-Verified is that the former runs an entire test file to evaluate the submitted patch, whereas we only run the contributed tests . Not running other test cases enables us to precisely compute coverage of . If the tests are relevant, they should cover the deleted lines in and the added lines in . We integrated the Python Coverage package into the 12 repositories and updated the test scripts to allow us to run specific test cases and compute coverage for them.
We define the metric that evaluates the quality of tests generated by a solution over a set of instances. It returns a number between 0 and 100, the higher the better. It is defined as 100 times the arithmetic mean of the per-instance scores:
Given a set of tests submitted for an instance, the per-instance score is a product of two factors:
The first factor is a binary correctness metric, using the indicator function for the tests failing on the old code times the indicator function for the tests passing on the new code. While the solution only has access to the old code , the evaluation metric also uses the hidden golden new code right after the issue was fixed.
The second factor is a fraction between 0 and 1 based on test coverage on the old and new code:
Adequacy focuses on just the coverage of lines added and deleted when going from the old code to the new code, because those are the most relevant lines to be tested. In the above, is the set of lines covered by running tests on code ; is the set of lines deleted by the PR patch; and is the set of lines added by the PR patch. We evaluate adequacy jointly for added and deleted lines, as some code patches may contain only added or deleted lines.